you are selling tickets to a football game. students cost $4 and general admission tickets cost $7. you sell 213 tickets ad collect $1146. how many of each type of ticket did you sell?
Answer by ewatrrr(24785)

 
Hi,
Let x represent number of general admission tickets
7x + 4(213-x) = 1146
7x + 852 - 4x = 1146
3x = 1146 - 852
3x = 294
x = 98, number of general admission tickets. 115 Student tickets 213-98
CHECKING our answer***
$7*98 + $4* 115 = $686 + $460 = $1146
Wish You the Best in your Studies.